# COBASE: A new copula-based shuffling method for ensemble weather forecast postprocessing

### v1

@@ 3.3 Existing parametric multivariate post-processing
The copula-based parametric correction steps for post-processing are summarized in Algorithm 2, which illustrates that any copula model can be used in place of the Gaussian copula. In our case studies, we estimate the copula parameters for standard parametric copula models using a 30-day rolling window, consistent with the EMOS procedure.
We notice that there has been limited research on the application of alternative copulas in the post-processing context, with notable exceptions such as Jobst et al. (2024) and Flos (2022). As highlighted in Table 1, while parametric copula-based methods exhibit considerable flexibility, a key limitation lies in their reliance on random sampling. Unlike structured techniques, such as the use of uniform quantiles, random draws provide a sparse and noisy representation of the univariate distributions. Since copula methods always generate corrected samples through random draws from a multivariate distribution, the resulting margins inherit this inefficiency. Consequently, univariate skill often suffers even when multivariate dependencies are well captured. In the next section, we provide evidence that this issue is substantial and would affect the performance of any multivariate parametric post-processing method.
@@
S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
Figure 1: Boxplots of Diebold-Mariano test statistics for CRPS of T2m, comparing EMOS-Q against EMOS-R (baseline). Results are presented for the Austrian and Dutch stations used in this study.
@@
3.3.1 Limitations of univariate random sampling in statistical post-processing
As mentioned earlier, the reliance of some univariate and multivariate post-processing techniques on random sampling can lead to poorly calibrated margins. To illustrate this, we compare the CRPS of T2m obtained for all stations in the study (Figure 1) using EMOS-R, i.e., random draws from the corrected Gaussian distribution, as the baseline. Figure 1 shows that drawing uniform quantiles instead (EMOS-Q) consistently outperforms EMOS-R across all stations.
We further examine how univariate random sampling affects multivariate corrections by comparing the Energy Score and the Variogram Score for the Sim Schaake method with different marginal samples, i.e., EMOS-R versus EMOS-Q. Essentially, the two methods compared in Figure 2 share the same dependence structure and only differ in the marginal sampling technique. The results show that Sim Schaake with randomly drawn margins (SimSSh-R) consistently performs worse than Sim Schaake with uniform quantile margins (SimSSh) in both scores across all groups, demonstrating that poor univariate sampling propagates into multivariate performance.
This investigation highlights how random sampling of marginal distributions can negatively affect multivariate corrections. Since copula-based techniques fundamentally rely on random draws from parametric multivariate distributions, this limitation may systematically compromise their performance. To overcome this limitation, we develop a new method, COBASE, which is presented in the following section.
@@ 3.4 A new copula-based shuffling method for ensemble postprocessing: COBASE
The proposed Copula-Based Shuffling of Ensemble (COBASE) method extends existing copula-based postprocessing techniques by introducing a shuffling of corrected univariate forecasts that replaces random marginal sampling. This approach preserves both realistic marginal behavior and the dependence structure defined by the chosen copula. The procedure is summarized in Algorithm 3. First, a parametric copula $C_{\alpha}$ is fitted to the data to capture the underlying dependence structure (Step S1). A random sample is then drawn from this copula to serve as the reference dependence structure $C_{ref}$ (Step S2). Next, standard univariate postprocessing, such as EMOS, is applied to each margin $l=1,\ldots,d$ to obtain corrected distributions $F_{l}$ (Step S3). Independent samples of size $N$ are drawn from these corrected margins using uniform quantiles (Step S4). This ensures that the marginal distributions are well-calibrated, including at the extremes, avoiding the shortcomings of random draws from a parametric multivariate distribution. Importantly, this step is flexible: any univariate sampling technique can be used, not necessarily uniform quantiles, for example, stratified sampling, allowing additional control over the sampled distributions. Finally, the univariate samples are shuffled according to the rank structure of $C_{ref}$ (Step S5). By doing so, COBASE preserves the dependence structure depicted by the parametric copula while leveraging the improved marginal calibration of advanced univariate sampling through, for example, uniform quantiles. As previously discussed, Table 1 provides an overview of the main multivariate post-processing methods and their key properties. In the table, COBASE stands out by combining the strengths of parametric and non-parametric two-step approaches: The parametric copula captures dependence, while uniform quantile sampling ensures well-calibrated margins. In the next section, we evaluate COBASE in two real-world case studies, comparing it with state-of-the-art methods to highlight its practical advantages. In the two case studies, we estimate the copula parameters for COBASE using a 30-day rolling window, consistent with the EMOS procedure and the standard copula models approaches.
@@
S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
@@
S3: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
S5: Shuffle the corrected univariate samples according to the ranks of $C_{ref}$.
Method Approach Dependence Free from long historical data Free from historical raw forecasts Flexible marginal sampling No restriction on the sample Size
Sim Schaake Shuffling Empirical copula (from past obs) × ×
Schaake shuffle Shuffling Empirical copula (from past obs) ×
ECC Shuffling Empirical copula (from raw forecasts) ×
Table 1: Summary of the multivariate post-processing methods analyzed in this work, alongside their main characteristics.
@@ 4 Results
We evaluate the proposed COBASE method in comparison with state-of-the-art post-processing approaches on the two real-world case studies described in Section 2. Although the focus of this work is not on univariate correction, CRPS results for 2-meter temperature across all stations are included in Table 3 for completeness. These results show that EMOS-Q, which uses uniform quantile sampling, consistently outperforms methods based on random sampling, clearly indicating that quantile-based approaches are generally preferable. The main focus of this section is on multivariate forecast performance, assessed using the Energy Score (ES) and the Variogram Score (VS). We begin by evaluating COBASE-GCA, the variant of COBASE using a Gaussian copula. COBASE-GCA is compared with Sim Schaake, Schaake Shuffle, ECC, and standard GCA.
@@ 5 Discussion and conclusion
This paper introduces COBASE, a new multivariate post-processing method that combines the structure of parametric copula models with the flexibility of non-parametric methods. By decoupling the sampling of margins from the copula-based dependence structure, COBASE addresses a key issue in the copula-based approaches for post-processing, namely, the degradation in forecast quality caused by random sampling from marginal distributions.
@@
Our results from two real-world case studies in Austria and the Netherlands show that COBASE consistently improves upon standard copula-based methods, including GCA, and outperforms Schaake Shuffle. Moreover, COBASE achieves performance comparable to widely used non-parametric techniques such as Sim Schaake and ECC, while avoiding their structural limitations and extensive data requirements.
A closer look at the scores in Table 4 reveals that COBASE-GCA performs similarly to Sim Schaake, even in the station groups where the latter shows a slight advantage. It is important to note, however, that the comparison is not entirely fair: Sim Schaake and Schaake Shuffle are trained on four years of data, whereas COBASE, like EMOS, uses only a 30-day rolling training window. This shorter training horizon makes COBASE substantially more flexible and easier to apply in operational contexts. For example, COBASE would be applicable to C-LAEF, while Sim Schaake would not be suitable in that case due to the lack of past raw forecasts. Additionally, COBASE shows robust performance even in relatively high-dimensional settings, such as the ECMWF ENS group with 12 variables, and no dimension-specific tuning.
The added value of COBASE becomes particularly evident in comparison to that of its corresponding standard copula-based variant. In this case, the performance gap is much more pronounced, confirming that the observed improvements stem from the structure of the COBASE method itself rather than from the specific copula family employed. This supports the main goal of the study: Not to identify the single best-performing copula, but to demonstrate that, across different configurations, COBASE consistently enhances the effectiveness of copula-based post-processing. In this sense, COBASE represents a more efficient, flexible, and broadly applicable alternative to conventional parametric multivariate approaches.
@@
Finally, while COBASE avoids the limitations of random marginal sampling, it still relies on random draws from the fitted copula to represent the multivariate structure. Future research could explore alternative multivariate sampling strategies to reduce sampling variability and better capture the characteristics of the dependence structure, especially in the tails of the multivariate distributions. These avenues offer promising directions for further strengthening COBASE’s performance across a wider range of postprocessing tasks.
To conclude, taken together, our findings demonstrate that COBASE already provides a simple, flexible, and effective framework for multivariate ensemble post-processing. By combining the strengths of parametric and non-parametric approaches, it improves forecast skill without requiring long historical records or raw ensemble data. The COBASE methodology is also particularly well suited for research on post-processing of multivariate dependence, as it allows a clear separation between the evaluation of the dependence structure and the marginals, thereby facilitating a better identification of potential improvements in the dependence component itself. These properties make COBASE particularly well-suited for operational settings with limited data or for newly developed forecasting systems, such as the C-LAEF system in Austria.
Building on this foundation, future work will focus on adjusting the method to higher-dimensional settings, improving copula selection strategies, and enhancing the multivariate sampling step. With these developments, COBASE has the potential to become a robust and scalable solution for a wide range of ensemble post-processing applications.
@@
The datasets analyzed during this study are subject to restrictions and are therefore not publicly available. To facilitate reproducibility, all analysis scripts and mock data that closely replicate the original datasets, allowing reproduction of the figures and tables, are available at https://github.com/elisaperrone/COBASE_github.
@@ Appendix A Abbreviations
Table 2: Overview of the statistical post-processing methods evaluated in the study, grouped by approach type, with their corresponding labels.

### v2

@@ 3.3 Existing parametric multivariate postprocessing
The copula-based parametric correction steps for postprocessing are summarized in Algorithm 2, which illustrates that any copula model can be used in place of the Gaussian copula. In both case studies, copula parameters for the parametric copula models are estimated using a rolling 30-day training window. In initial experiments, longer training windows (e.g., 50 days) were also considered but did not lead to clear performance changes. Using longer windows increases the risk of mixing different seasonal statistics in the training data, while overly short windows may adversely affect the quality of dependence estimation. We therefore adopt a rolling 30-day window to ensure consistency with the EMOS-based univariate postprocessing, so that both univariate and multivariate corrections rely on the same pool of past observations. In our study, copula parameters are estimated on each rolling training window using pseudo-observations obtained from the marginal CDF values. For the Gaussian and Archimedean copulas considered here, the copula parameters are fitted using inversion of Kendall’s tau (Hofert et al., 2018), which provides a fast and robust moment-based estimator.
Although parametric copula-based approaches also rely on past observations for parameter estimation, their dependence on historical data is less pronounced than for empirical shuffling methods such as the Schaake Shuffle or Sim Schaake, which directly use past observations as dependence templates. In parametric copula models, the estimated dependence structure can deviate from any individual historical realization, allowing for greater flexibility in representing multivariate dependence.
We notice that there has been limited research on the application of alternative copulas in the postprocessing context, with notable exceptions such as Jobst et al. (2024) and Flos (2022). As highlighted in Table 1, while parametric copula-based methods exhibit considerable flexibility, a key limitation lies in their reliance on random sampling. Unlike structured techniques, such as the use of uniform quantiles, random draws provide a sparse and noisy representation of the univariate distributions. Since copula methods always generate corrected samples through random draws from a multivariate distribution, the resulting margins inherit this inefficiency. Consequently, univariate skill often suffers even when multivariate dependencies are well captured. In the next section, we provide evidence that this issue is substantial and would affect the performance of any multivariate parametric postprocessing method.
@@
Step1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
Step2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
Figure 1: Boxplots of Diebold-Mariano test statistics Diebold and Mariano (1995) for CRPS of T2m, comparing EMOS-Q against EMOS-R (baseline). Results are presented for the Austrian and Dutch stations used in this study.
@@
3.3.1 Limitations of univariate random sampling in statistical postprocessing
As mentioned earlier, the reliance of some univariate and multivariate postprocessing techniques on random sampling can lead to poorly calibrated margins. To illustrate this, we compare the CRPS of T2m obtained for all stations in the study (Figure 1) using EMOS-R, i.e., random draws from the corrected Gaussian distribution, as the baseline. Figure 1 shows that drawing uniform quantiles instead (EMOS-Q) consistently outperforms EMOS-R across all stations.
We further examine how univariate random sampling may affect multivariate corrections. To illustrate this effect, we focus on the Sim Schaake method, which has been shown to perform best for the ALADIN-LAEF system (Perrone et al., 2020) and remains competitive for the ECMWF dataset considered here (Schefzik, 2016). We compare the Energy Score and the Variogram Score for Sim Schaake using different marginal sampling strategies, namely EMOS-R and EMOS-Q. Essentially, the two methods compared in Figure 2 share the same dependence structure and only differ in the marginal sampling technique. The results show that Sim Schaake with randomly drawn margins (SimSSh-R) consistently performs worse than Sim Schaake with uniform quantile margins (SimSSh) in both scores across all groups, demonstrating that poor univariate sampling propagates into multivariate performance.
This investigation highlights how random sampling of marginal distributions can negatively affect multivariate corrections. Since copula-based techniques fundamentally rely on random draws from parametric multivariate distributions, this limitation may negatively affect their performance. To overcome this limitation, we develop a new method, COBASE, which is presented in the following section.
@@ 3.4 A new copula-based shuffling method for ensemble postprocessing: COBASE
The proposed Copula-Based Shuffling of Ensemble (COBASE) method extends existing copula-based postprocessing techniques by introducing a shuffling of corrected univariate forecasts that replaces random marginal sampling. This approach preserves both realistic marginal behavior and the dependence structure defined by the chosen copula. The procedure is summarized in Algorithm 3. First, a parametric copula $C_{\alpha}$ is fitted to the data to capture the underlying dependence structure (Step S1). A random sample is then drawn from this copula to serve as the reference dependence structure $C_{ref}$ (Step S2). Next, standard univariate postprocessing, such as EMOS, is applied to each margin $l=1,\ldots,d$ to obtain corrected distributions $F_{l}$ (Step3). Independent samples of size $N$ are drawn from these corrected margins using uniform quantiles (Step S4). This ensures that the marginal distributions are well-calibrated, including at the extremes, avoiding the shortcomings of random draws from a parametric multivariate distribution. Importantly, this step is flexible: any univariate sampling technique can be used, not necessarily uniform quantiles, for example, stratified sampling, allowing additional control over the sampled distributions. Finally, the univariate samples are shuffled according to the rank structure of $C_{ref}$ (Step S5). By doing so, COBASE preserves the dependence structure depicted by the parametric copula while leveraging the improved marginal calibration of advanced univariate sampling through, for example, uniform quantiles. As previously discussed, Table 1 provides an overview of the main multivariate postprocessing methods and their key properties. In the table, COBASE stands out by combining the strengths of parametric and nonparametric two-step approaches: The parametric copula captures dependence, while uniform quantile sampling ensures well-calibrated margins. In the next section, we evaluate COBASE in two real-world case studies, comparing it with state-of-the-art methods to highlight its practical advantages. In the two case studies, we estimate the copula parameters for COBASE using a 30-day rolling window, consistent with the EMOS procedure and the standard copula models approaches.
@@
Step1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
@@
Step3: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
Step5: Shuffle the corrected univariate samples according to the ranks of $C_{ref}$.
@@ 4 Results
We evaluate the proposed COBASE method in comparison with state-of-the-art postprocessing approaches on the two real-world case studies described in Section 2. Although the focus of this work is not on univariate correction, CRPS results for 2-meter temperature across all stations are included in Table 3 for completeness. These results show that EMOS-Q, which uses uniform quantile sampling, outperforms methods based on random sampling on average, indicating that quantile-based approaches are generally preferable in this setting. The reported differences are based on mean CRPS values computed over the full evaluation period. Although averaging may mask the influence of individual validation days, the overall results consistently favor quantile-based sampling. Small differences in mean CRPS among methods relying on random sampling of the marginal distributions may partly reflect the variability induced by the random draws, which is a known limitation of EMOS-R and traditional copula-based approaches. This source of variability is avoided by the deterministic sampling method used in EMOS-Q. The main focus of this section is on multivariate forecast performance, assessed using the Energy Score (ES) and the Variogram Score (VS). We begin by evaluating COBASE-GCA, the variant of COBASE using a Gaussian copula. COBASE-GCA is compared with Sim Schaake, Schaake Shuffle, ECC, and standard GCA.
@@ 5 Discussion and conclusion
This paper introduces COBASE, a new multivariate postprocessing method that combines the structure of parametric copula models with the flexibility of nonparametric methods. By decoupling the sampling of margins from the copula-based dependence structure, COBASE addresses a key issue in the copula-based approaches for postprocessing, namely, the degradation in forecast quality caused by random sampling from marginal distributions.
@@
Our results from two real-world case studies in Austria and the Netherlands show that COBASE consistently improves upon standard copula-based methods, including GCA, and outperforms Schaake Shuffle. Moreover, COBASE achieves performance comparable to widely used nonparametric techniques such as Sim Schaake and ECC, while avoiding their structural limitations and extensive data requirements.
A closer look at the scores in Table 4 reveals that COBASE-GCA performs similarly to Sim Schaake, even in the station groups where the latter shows a slight advantage. It is important to note, however, that the comparison is not entirely fair: Sim Schaake and Schaake Shuffle are trained on four years of data, whereas COBASE, like EMOS, uses only a 30-day rolling training window. This shorter training horizon makes COBASE substantially more flexible and easier to apply in operational contexts. For example, COBASE would be applicable to C-LAEF, whereas Sim Schaake would not be suitable in that case due to the lack of past raw forecasts. While Lakatos et al. (2023) report that increasing the number of historical pool dates for the Schaake Shuffle does not necessarily lead to improved forecast performance, the Schaake Shuffle still relies on historical dependence templates and exhibits weaker average performance in our experiments. In contrast, COBASE achieves improved skill on average while requiring only a short rolling training window, making it a more practical choice in operational settings. Additionally, COBASE shows robust performance even in relatively high-dimensional settings, such as the ECMWF ENS group with 12 variables, and no dimension-specific tuning. Another important aspect is that COBASE, as any parametric copula-based method, can be applied to both exchangeable and nonexchangeable ensemble systems, whereas this may not necessarily hold for ECC and Sim Schaake, depending on the underlying assumptions and, in the latter case, on the chosen similarity criterion. While ECC formally relies on ensemble exchangeability, previous work in Perrone et al. (2020) has shown that nonexchangeability does not necessarily prevent its application, although it requires careful diagnostic analysis to identify potential anomalous dependence structures in the raw forecasts. Such considerations depend on the specific characteristics of the NWP system and are therefore model-dependent. For this reason, we include ECC in this study for comparison purposes.
The added value of COBASE becomes particularly evident in comparison to that of its corresponding standard copula-based variant. In this case, the performance gap is much more pronounced, confirming that the observed improvements stem from the structure of the COBASE method itself rather than from the specific copula family employed. This supports the main goal of the study: Not to identify the single best-performing copula, but to demonstrate that, across different configurations, COBASE consistently enhances the effectiveness of copula-based postprocessing. In this sense, COBASE represents a more efficient, flexible, and broadly applicable alternative to conventional parametric multivariate approaches.
@@
Then, while COBASE avoids the limitations of random marginal sampling, it still relies on random draws from the fitted copula to represent the multivariate structure. Future research could explore alternative multivariate sampling strategies to reduce sampling variability and better capture the characteristics of the dependence structure, especially in the tails of the multivariate distributions. These avenues offer promising directions for further strengthening COBASE’s performance across a wider range of postprocessing tasks.
Finally, an important direction for future work is the development of controlled simulation studies to further assess the performance of COBASE relative to benchmark methods, and to systematically investigate the role of ensemble size and ensemble exchangeability.
To conclude, taken together, our findings demonstrate that COBASE already provides a simple, flexible, and effective framework for multivariate ensemble postprocessing. By combining the strengths of parametric and nonparametric approaches, it improves forecast skill without requiring long historical records or raw ensemble data. The COBASE methodology is also particularly well suited for research on postprocessing of multivariate dependence, as it allows a clear separation between the evaluation of the dependence structure and the marginals, thereby facilitating a better identification of potential improvements in the dependence component itself. These properties make COBASE particularly well-suited for operational settings with limited data or for newly developed forecasting systems, such as the C-LAEF system in Austria.
Building on this foundation, future work will focus on adjusting the method to higher-dimensional settings, improving copula selection strategies, and enhancing the multivariate sampling step. With these developments, COBASE has the potential to become a robust and scalable solution for a wide range of ensemble postprocessing applications.
Acknowledgment
We thank the two anonymous reviewers and the associate editor for their thoughtful comments, which helped improve an earlier version of the manuscript.
Author Contribution Indication
Maurits Flos: Conceptualization, Data curation, Formal analysis, Investigation, Software, Validation
Bastien François: Conceptualization, Data curation, Investigation, Validation, Visualization, Writing – original draft, Writing – review & editing
Irene Schicker: Data curation, Writing – original draft, Writing – review & editing
Kirien Whan: Conceptualization, Data curation, Investigation, Software, Validation, Visualization, Writing – original draft, Writing – review & editing
Elisa Perrone: Conceptualization, Formal analysis, Investigation, Methodology, Software, Supervision, Validation, Visualization, Writing – original draft, Writing – review & editing
@@ Acknowledgment
The data that support the findings of this study are available from the corresponding author upon reasonable request. To facilitate reproducibility, all analysis scripts and mock data that closely replicate the original datasets, allowing reproduction of the figures and tables, are available at https://github.com/elisaperrone/COBASE_github.
Conflict of interest statement
The authors have no conflicts of interest to declare.
@@ Appendix A Abbreviations
Table 2: Overview of the statistical postprocessing methods evaluated in the study, grouped by approach type, with their corresponding labels.